# PARCO: Learning Parallel Autoregressive Policies for Efficient Multi-Agent Combinatorial Optimization

### v1

@@ 5.3 Experimental Results
Main experiments
The main experiments showcasing the performance of PARCO are in Table 1, Table 2, Table 3 for HCVRP, OMDCPDP, and FFSP, respectively. Our PARCO consistently outperforms SotA neural baselines across a variety of experiments. We additionally note another property of PARCO: unlike neural baselines in HCVRP and MAPDP in the OMDCPDP, which are trained specifically for a single size and number of agents, PARCO is a single model trained on multiple location and agent distributions at the same time; nonetheless, our single model can outperform baselines trained ad hoc on specific distributions, demonstrating our method’s flexibility.
Generalization
We additionally study the generalization performance in Table 2 on the right, which are the sizes and number of agents unseen during training in OMDCPDP. Unlike MAPDP, which is constrained to a specific number of agents, PARCO can successfully generalize to unseen sizes and $m$. Notably, PARCO can even outperform Google ORTools for larger-scale instances, demonstrating remarkable scalability.
Table 4: Effect of different ##\## of agents in FFSP.
$J\times S\times m$ Metric MatNet PARCO
$50\times 3\times 6$ Obj. 35.69 33.58
Gap 6.3% 0.0%
Duration 1.1h 10.3m
Avg. Steps 200.05 19.10
$50\times 3\times 8$ Obj. 28.59 25.62
Gap 11.6% 0.0%
Duration 1.5h 9.4m
Avg. Steps 205.95 16.80
$50\times 3\times 10$ Obj. 25.05 21.46
Gap 16.7% 0.0%
Duration 1.8h 7.5m
Avg. Steps 215.85 13.45
@@
We showcase the importance of Communication Layers in Fig. 3 (a). We benchmark 1) No Communication (only context features), 2) communication via an MLP, 3) communication via an MHA layer, and 4) Our Communication layer. Our Communication Layer consistently outperforms other methods.
@@
Fig. 3 (b) shows the effect of the Priority-based Conflict Handler with different priorities methods $\bm{p}$ in the OMDCPDP. We consider the following: 1) Random: the priority is chosen randomly as in Zong et al. (2022), 2) First: the priority is chosen by the agent index $k=1,\dots,m$, 3) Smallest: gives priority to the agent that has traveled the least, 4) Closest: priority is given to the closest agent and 5) High Probability: priority is given to the agent whose probability of selecting the conflicting action is the highest. High Probability can consistently outperform other methods. We note that this is also the most general and problem-agnostic since methods 2-4 are problem-specific.
Scalability
Finally, we showcase PARCO’s scalability in terms of speed in large-scale instances in Fig. 3 (c) compared to the constructive autoregressive ET (Son et al., 2024). Interestingly, PARCO’s gap grows with more agents. This is because our method can fully exploit agent parallelism and can thus be a strong candidate for large-scale real-time CO applications.
@@ 6 Conclusion
In this paper, we introduced a novel approach, PARCO - Parallel AutoRegressive Combinatorial Optimization - that utilizes parallel autoregressive (AR) policies to tackle multi-agent NCO problems. We introduced a new Communication Layer to allow for the agents to effectively coordinate their next steps during decoding, alongside a Multiple Pointer Mechanism coupled with a Priority-based Conflicts Handler that can generate feasible solutions efficiently. We validate the effectiveness of PARCO through extensive experiments with different CO problems from the routing and scheduling domain and demonstrate its competitive performance against classical heuristic and neural baselines.
While PARCO is already faster than most evaluated neural baselines and decreases the number of required decoding steps by more than 90% in FFSP50 instances with 10 agents as shown in Table 4, it still does not achieve the potential reduction from $\mathcal{O}(N)$ to $\mathcal{O}(\frac{N}{m})$ in many cases. This is mainly attributed to the fact that agents can have conflicts that we currently resolve with the priority-based conflict handler. Giving one agent precedence and forcing others to stay in their current position results in more than necessary decoding steps. In future work, we will try to address this limitation by constructing a custom loss function, which penalizes conflicts and makes agents learn to avoid them during training. Further, we plan to extend our work with learnable improvement methods to enable fast first solutions coupled with powerful local search. Given PARCO’s solution quality and efficiency, a further exciting avenue of future research is to couple our model with population-based approaches (Grinsztajn et al., 2022; Chalumeau et al., 2023; Hottung et al., 2024) that could search the latent space for diverse and better solutions.
@@ Acknowledgements
We gratefully acknowledge OMELET for supporting us with computing. We also thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on.
Appendix
[…]
Appendix A Problem Descriptions
@@ A.3 FFSP
A.3.1 Problem Definition
The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed across $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The FFSP can naturally be viewed as a multi-agent CO problem by considering each machine as an agent that constructs its own schedule. Adhering to autoregressive CO, agents construct the schedule sequentially, selecting one job (or no job) at a time. The job selected by a machine (agent) at a specific stage in the decoding process is scheduled at the earliest possible time, that is, the maximum of the time the job becomes available in the respective stage (i.e., the time the job finished on prior stages) and the machine becoming idle. The process repeats until all jobs for each stage have been scheduled, and the ultimate goal is to minimize the makespan, i.e., the total time required to complete all jobs.
A.3.2 Mathematical Formulation
@@
$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk}-1),0\right)$ $i=1,\dots,S$ (32)
@@
Constraint Explanations: Here, the objective function Eq. 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set Eq. 30 ensures that each job is assigned to exactly one machine at each stage. Constraint sets Eq. 31 through Eq. 35 define the precedence relationships between jobs within a stage. Specifically, constraint set Eq. 31 ensures that a job has no precedence relationship with itself. Constraint set Eq. 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set Eq. 33 dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets Eq. 34 and Eq. 35 restrict a job to having at most one preceding job and one following job.
Moving on, constraint set Eq. 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set Eq. 37. Finally, constraint set Eq. 38 ensures that no more than one job can be processed on the same machine simultaneously.
@@ B.1.1 Baselines
The Slack Induction by String Removals (SISR) approach (Christiaens and Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
@@
The Genetic Algorithm (GA) (Karakatič and Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. Particularly effective in Multi-Depot Vehicle Routing Problems (MDVRP), GA quickly generates adequate solutions with reasonable computational resources. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
@@
The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from Liu et al. (2024c) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
@@
The Attention Model (AM) (Kool et al., 2018) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
@@
The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP, and integrate vehicle features into both the input layer and the context encoder.
@@
The DRL approach for solving HCVRP by Li et al. (2022) employs a transformer architecture similar to Kool et al. (2018) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents $m$, each with always the same characteristics.
@@
The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024c) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to changes in vehicle and customer numbers, ensuring robust performance across different scenarios. For our study, 2D-Ptr was configured following the original study’s parameters for consistent comparison with 8 heads, 128 hidden dimensions for model structure; $\mathrm{lr}=10^{-4},\lambda_{\mathrm{Adam}}=0.995,\kappa_{\mathrm{L2}}=3$ for turning; 50 epochs, 2500 batches per epoch, 512 instances per batch for training.
@@ B.1.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents $m$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in the Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $k$, where $k=1,\ldots,m$, is assigned a capacity $Q_{k}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{k}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
@@
Testing is performed on the 1280 instances per test setting from Liu et al. (2024c). In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward). We report optimality gaps to ORTools and solution time in seconds in parentheses.

### v2

@@ 5.2 Experimental Results
Table 2: Large-scale generalization for unseen numbers of nodes $N$ and agents $M$ (up to 10×10$\times$10 × of those seen during training).
$N$ 500 1000
$M$ 50 75 100 100 150 200
Metric Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time Obj. Gap Time
OR-Tools 290.79 7.81% 300s 223.72 4.68% 300s 192.59 1.51% 300s 780.89 33.93% 600s 642.54 36.47% 600s 584.27 36.81% 600s
HAM (g.) 410.95 55.44% 1.03s 310.17 48.72% 1.13s 204.00 9.91% 1.23s 710.64 40.92% 2.43s 554.15 39.41% 2.66s 388.56 8.88% 2.91s
PARCO (g.) 268.56 1.41% 0.02s 211.21 1.13% 0.02s 187.37 0.83% 0.01s 510.61 0.81% 0.02s 401.46 0.64% 0.02s 359.98 0.49% 0.02s
HAM (s.) 409.67 54.95% 1.19s 305.16 46.32% 1.31s 203.50 9.64% 1.39s 708.55 40.50% 2.94s 552.76 39.06% 3.22s 384.10 7.63% 3.46s
PARCO (s.) 264.38 0.00% 0.03s 208.56 0.00% 0.02s 185.61 0.00% 0.02s 504.30 0.00% 0.79s 397.51 0.00% 0.91s 356.87 0.00% 1.26s
We present the main empirical results for HCVRP, OMDCPDP, and FFSP in Table 1, reporting the average objective function values (Obj.), gaps to the best-known solutions, and inference times for solving each single instance of each problem. For neural baselines, we evaluate both greedy (g.) and sampling (s.) performance, using 1280 sampled solutions for routing problems and 128 for FFSP.
In HCVRP, PARCO outperforms all neural baselines in solution quality and speed while providing solutions at a fraction of the solving time required by traditional solvers. In OMDCPDP, our model surpasses all baselines, including OR-Tools. Notably, while the AR baseline HAM struggles with a larger number of agents $M$, PARCO’s parallel AR method maintains strong performance across all scales. In FFSP, PARCO outperforms traditional solvers (e.g., Gurobi cannot find solutions in time for $N>20$), dispatching rules, and MatNet in all tested scenarios while being more than 4×4$\times$4 × faster. Furthermore, similar to our results in routing, PARCO’s advantage becomes even more pronounced in instances with a larger number of agents where PARCO generates higher-quality schedules through effective agent coordination at a fraction of the cost of MatNet.
5.3 Analysis
@@
We showcase the importance of Communication Layers in Figure 3(a). We benchmark different ways to modify decoder queries (see Section 4.5) with 1) No communication (only context features), 2) MLP, 3) MHA, 4) our transformer-based Communication Layers. Our Communication Layers consistently outperform other methods. Figure 3(b) further shows the effect of communication, which can not only improve performance but also reduce training time by reducing the number of steps, thereby offsetting the additional latency introduced by the Communication Layers.
@@
We study the effect of conflict handlers in Figure 3(c). We compare 1) the random handler from MAPDP and our proposed Priority-based Conflict Handlers in different configurations, namely with priorities based on heuristics, including 2) “smallest” prioritizing the agent with the lowest cost so far, and 3) “closest”, prioritizing agents closer to the corresponding node, and finally 4) “learned” with learned priorities (see Section 4.6). PARCO’s Priority-based Conflict Handlers with learned priorities consistently outperform other methods, which also enjoy a relative reduction in the number of steps for constructing a solution, e.g., with a $4\%$ reduction in conflict rates.
Large-Scale Generalization
We study the zero-shot large-scale generalization performance of PARCO on OMDCPDP and report the results in Table 2 for out-of-distribution numbers of nodes $N$ and agents $M$, both up to 10×10$\times$10 × those seen in training. The autoregressive HAM baseline cannot generalize well to such scales due to the lack of communication and robust parallel construction. We note that MAPDP cannot be applied to an unseen number of agents because of its inflexible decoder structure. Conversely, our method outperforms all baselines, even including OR-Tools with a 10-minute solving time per instance for $N=1000$, making PARCO a powerful candidate for real-time deployment.
Figure 4: PARCO vs AR inference time. PARCO can construct solutions faster with more agents $M$ thanks to parallel decoding.
PARCO vs AR Models Scalability
Finally, we showcase PARCO’s scalability against autoregressive methods in Figure 4. Notably, compared to AR models (e.g., DPN, HAM), PARCO achieves significantly faster solution speeds, with efficiency improving as the number of agents $M$ increases. Thanks to parallel decoding, fewer solution construction steps for larger $M$ lead to substantially lower latency.
@@ 6 Conclusion
We introduced PARCO, a learning model to tackle multi-agent combinatorial optimization problems efficiently via parallel autoregressive solution construction. By integrating transformer-based Communication Layers, a Multiple Pointer Mechanism, and Priority-based Conflict Resolution, PARCO enables effective agent coordination and significantly reduces computational latency. Our extensive experiments on multi-agent vehicle routing and scheduling demonstrate that PARCO consistently outperforms state-of-the-art learning-based solvers, achieving higher solution quality while maintaining superior efficiency.
Limitations & Future Work
Although PARCO can efficiently solve multi-agent CO problems with a defined number of agents $M$, it cannot be directly applied to CO tasks where solutions are constructed via an unspecified $M$. In future work, we plan to explore multi-agent CO problems with an unspecified number of agents, which could be achieved by either rolling out a batch of several values of $M$ until an optimal solution is reached or by employing a prediction module to predict an optimal agent number.
@@ Acknowledgements
We thank people in the AI4CO open research community who have contributed to RL4CO, which this work is based on. This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT)[2022-0-01032, Development of Collective Collaboration Intelligence Framework for Internet of Autonomous Things].
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Problem Definitions
@@ A.3 FFSP
The flexible flow shop problem (FFSP) is a challenging and extensively studied optimization problem in production scheduling, involving $N$ jobs that must be processed by a total of $M$ machines divided into $i=1\dots S$ stages, each with multiple machines ($m_{i}>1$). Jobs follow a specified sequence through these stages, but within each stage, any available machine can process the job, with the key constraint that no machine can handle more than one job simultaneously. The FFSP can naturally be viewed as a multi-agent CO problem by considering each machine as an agent that constructs its own schedule. Adhering to autoregressive CO, agents construct the schedule sequentially, selecting one job (or no job) at a time. The job selected by a machine (agent) at a specific stage in the decoding process is scheduled at the earliest possible time, that is, the maximum of the time the job becomes available in the respective stage (i.e., the time the job finished on prior stages) and the machine becoming idle. The process repeats until all jobs for each stage have been scheduled, and the ultimate goal is to minimize the makespan, i.e., the total time required to complete all jobs.
Mathematical Model
@@
$\sum_{j=1}^{N}\sum_{l=1}^{N}Y_{ilj}$ $=\sum_{k=1}^{m_{i}}\max\left(\sum_{j=1}^{n}(X_{ijk})-1,0\right)$ $i=1,\dots,S$ (32)
@@
Constraint Explanations: Here, the objective function Equation 29 minimizes the makespan of the resulting schedule, that is, the completetion time of the job that finishes last. The schedule has to adhere to several constraints: First, constraint set Equation 30 ensures that each job is assigned to exactly one machine at each stage. Constraint sets Equation 31 through Equation 35 define the precedence relationships between jobs within a stage. Specifically, constraint set Equation 31 ensures that a job has no precedence relationship with itself. Constraint set Equation 32 ensures that the total number of precedence relationships in a stage equals $N-m_{i}$ minus the number of machines with no jobs assigned. Constraint set Equation 33 dictates that precedence relationships can only exist among jobs assigned to the same machine. Additionally, constraint sets Equation 34 and Equation 35 restrict a job to having at most one preceding job and one following job.
Moving on, constraint set Equation 36 specifies that the completion time of a job in the first stage must be at least as long as its processing time in that stage. The relationship between the completion times of a job in consecutive stages is described by constraint set Equation 37. Finally, constraint set Equation 38 ensures that no more than one job can be processed on the same machine simultaneously.
@@ B.1.1 Baselines
We follow the experimental setup of Liu et al. (2024b) for baselines, with additional baselines hyperparameter details reported in their respective papers.
@@
The Slack Induction by String Removals (SISR) approach (Christiaens & Vanden Berghe, 2020) offers a heuristic method for addressing vehicle routing problems (VRPs), focusing on simplifying the optimization process. It combines techniques for route dismantling and reconstruction, along with vehicle fleet minimization strategies. SISR is applied across various VRP scenarios, including those with specific pickup and delivery tasks. In our experiments, we adhere to the hyperparameters provided in the original paper with $\bar{c}=10,L^{\mathrm{max}}=10,\alpha=10^{-3},\beta=10^{-2},T_{0}=100,T_{f}=1,\mathrm{iter}=3\times 10^{5}\times N$.
@@
The Genetic Algorithm (GA) (Karakatič & Podgorelec, 2015) is used to address vehicle routing problems (VRPs) and other NP-hard challenges by simulating natural evolutionary processes. GA generates adequate solutions with reasonable computational resources. Our experiment follows the same carefully tuned hyperparameters from (Liu et al., 2024b) with $n=200,\mathrm{iter}=40\times N,P_{m}=0.8,P_{c}=1$.
@@
The Simulated Annealing (SA) method (İlhan, 2021) targets the capacitated vehicle routing problem (CVRP) using a population-based approach combined with crossover operators. It incorporates local search and the improved 2-opt algorithm to refine routes alongside crossover techniques to speed up convergence. In our experiment, we follow the same carefully tuned hyperparameters from (Liu et al., 2024b) with $T_{0}=100,T_{f}=10^{-7},L=20\times N,\alpha=0.98$.
@@
The Attention Model (AM) (Kool et al., 2019a) applies the attention mechanism to tackle combinatorial optimization problems like the Traveling Salesman and Vehicle Routing Problems. It utilizes attention layers for model improvement and trains using REINFORCE with deterministic rollouts. In our studies, we adopt adjustments from the DRLLi framework, which involves selecting vehicles sequentially and then choosing the next node for each. Additionally, vehicle-specific features are incorporated into the context vector generation to distinguish between different vehicles.
@@
The Equity-Transformer (ET) approach (Son et al., 2024) addresses large-scale min-max routing problems by employing a sequential planning approach with sequence generators like the Transformer. It focuses on equitable workload distribution among multiple agents, applying this strategy to challenges like the min-max multi-agent traveling salesman and pickup and delivery problems. In our experiments, we modify the decoder mask in ET to generate feasible solutions for HCVRP and integrate vehicle features into both the input layer and the context encoder, similarly to the setting of Liu et al. (2024b).
DPN
The Decoupling-Partition-Navigation (DPN) approach (Zheng et al., 2024a) is a SOTA sequential planning AR baseline that tackles min-max vehicle routing problems (min-max VRPs) by explicitly separating the tasks of customer partitioning and route navigation. It introduces a Partition-and-Navigation (P&N) Encoder to learn distinct embeddings for these tasks, an Agent-Permutation-Symmetric (APS) loss to leverage routing symmetries, and a Rotation-Based Positional Encoding to enhance generalization across different depot locations. We employ a similar setting as ET.
@@
The DRL approach for solving HCVRP by Li et al. (2022) employs a transformer architecture similar to Kool et al. (2019a) in which the vehicle and node selection happens in two steps via a two selection decoder, thus requiring two actions. We employ their original model with additional context of variable vehicle speeds, noting that in the original setting each model was trained on a single distribution of number of agents ${M}$, each with always the same characteristics.
@@
The 2D Array Pointer network (2D-Ptr) (Liu et al., 2024b) addresses the heterogeneous capacitated vehicle routing problem (HCVRP) by using a dual-encoder setup to map vehicles and customer nodes effectively. This approach facilitates dynamic, real-time decision-making for route optimization. Its decoder employs a 2D array pointer for action selection, prioritizing actions over vehicles. The model is designed to adapt to vehicle and customer numbers changes, ensuring robust performance across different scenarios.
@@ B.1.2 Datasets
Neural baselines were trained with the specific number of nodes $N$ and number of agents ${M}$ they were tested on. In PARCO, we select a varying size and number of customer training schemes: at each training step, we sample $N\sim\mathcal{U}(60,100)$ and $m\sim\mathcal{U}(3,7)$. As we show in Table 1, a single PARCO model can outperform baseline models even when they were fitted on a specific distribution. The coordinates of each customer location $(x_{i},y_{i})$, where $i=1,\ldots,N$, are sampled from a uniform distribution $\mathcal{U}(0.0,1.0)$ within a two-dimensional space. The depot location is similarly sampled using the same uniform distribution. The demand $d_{i}$ for each customer $i$ is also drawn from a uniform distribution $\mathcal{U}(1,10)$, with the depot having no demand, i.e., $d_{0}=0$. Each vehicle $m$, where $m=1,\ldots,{M}$, is assigned a capacity $Q_{m}$ sampled from a uniform distribution $\mathcal{U}(20,41)$. The speed $f_{m}$ of each vehicle is uniformly distributed within the range $\mathcal{U}(0.5,1.0)$.
@@
Testing is performed on the 1280 instances per $(N,M)$ test setting from Liu et al. (2024b). In Table 1, (g.) refers to the greedy performance of the model, i.e., taking a single trajectory by taking the maximum action probability; (s.) refers to sampling $1280$ solutions in the latent space and selecting the one with the lowest cost (i.e., highest reward).